A student taking a quiz finds on a reference sheet the two equations f = 1 / T and =square root of (1/T).
She has forgotten what T represents in each question.
Use dimensional analysis to determine the units required for T in each question.
Explain how you can identify the physical quantity each T represents from the units.
1
Expert's answer
2020-10-01T10:11:29-0400

solution

given formula


"f=\frac{1}{T}"


f is the frequency so unit of frequency is "s^{-1}"

by using dimension analysis

"T=\frac{1}{f}"


so

"[ \quad T\quad]=\frac{1}{s^{-1}}"


=s.

so unit of T is second and T is time.


"v=\sqrt \frac{T}{\mu}"


and v is velocity and mu is linear density

"T=v^2\mu"

"[\quad T\quad]=[\quad v^2\mu \quad]"


"[\quad T\quad]=m^2/s^2\cdot kg/m\\"


"[\quad T\quad]=\frac{kg\cdot m}{s^2}=N"

there T represents force.